OpenStudy (anonymous):

A xylophone has a set of vibrating bars. A bar with a frequency of 343 Hz is struck. A very short time later, a second bar is also sounded, and one hears the volume of the xylophone increase and decrease 4 times each second (called beats). What are the possible frequencies of the second bar (in Hz)? (Give all correct answers, i.e. C, AE, ABCD, etc...) A) 335 B) 348 C) 347 D) 351 E) 339

OpenStudy (anonymous):

I can't remember the equations here, but we want the peaks of the two waves to be aligned every quarter-second.

OpenStudy (anonymous):

:S i am unsure as well, we havent leant this. hold on may this help

OpenStudy (anonymous):

s(t)=[2scosw't]coswt

OpenStudy (anonymous):

does that help? s1=scoswt s2=scoswt thats the resultant

OpenStudy (jamesj):

Right, this equation comes from adding two waves together: \[ \cos(f_1.t) + \cos(f_2.t) = 2 \cos(w_1.t).\cos(w_2.t) \] where \[ w_1 = \frac{f_1 + f_2}{2}, \ w_2 = \frac{f_1 - f_2}{2} \] Now, if you're hearing beats, what you're hearing is the beats from the slower of these two frequencies, i..e, from \( w_2 \). Given that, you can now solve for the frequency of the other sound.

OpenStudy (anonymous):

i still dont get it :/

OpenStudy (jamesj):

Here's an example of beat behavior. http://www.wolframalpha.com/input/?i=cos%28440t%29+%2B+cos%28445t%29

OpenStudy (jamesj):

The point is, using superposition, if you have two sounds of a single frequency, then the sound you hear is just the sum of them: for example cos(f1.t) + cos(f2.t)

OpenStudy (jamesj):

That's the sound you would hear if you were hearing two sounds of equal amplitude, one of frequency f1, one of frequency f2. So far so good?

OpenStudy (anonymous):

ya

OpenStudy (jamesj):

Now, if f1 and f2 are close, then you will get "beats", as the two sounds interfere with each other. The beats can be calculated as I've shown above algebraically. The sound you hear as the beat sound, is the larger envelope of oscillation, the one with the lower frequency, the \[ \cos\left( \frac{f_1 - f_2}{2}t \right) \] term. And in inside that larger oscillation is the othe other term, the cos[(f1 + f2)/2.t] term frantically oscillating up and down. But

OpenStudy (jamesj):

But what your ear is hearing as 'beats' is the longer oscillation, as the sound diminishes to zero, then peaks, then zero, then peaks, etc.

OpenStudy (jamesj):

In your problem, your are hearing those oscillations 4 times a second. That means that \[ \frac{f_1 - f_2}{2} \] is such that you'll hear four beats. What must that number be equal to have four beats, or equivalently, four "zero" points?

OpenStudy (anonymous):

i get what you mean, but you only have 1 frequency, i dont get the +/- 4/second

OpenStudy (jamesj):

You have the original sound which has a frequency of f_1 = 343 Hz. Then a second sound is started with a frequency of f_2. Now you hear beats. So the question is: what values of f_2 make it possible for you to hear 4 beats/second?

OpenStudy (jamesj):

So your ear is hearing cos(f1.t) + cos(f2.t)

OpenStudy (anonymous):

wouldnt it b then ACDE?

OpenStudy (jamesj):

What value must (f1 - f2)/2 have?

OpenStudy (anonymous):

Since they will produce a difference of 4=|343-347|?

OpenStudy (jamesj):

No. What you need is you need the function cos[(f1 - f2)/2.t] to have four beats, which is equivalent it having 4 zeros every second. What value or values of (f1-f2)/2 make that possible?

OpenStudy (anonymous):

OH its D and A no?

OpenStudy (jamesj):

What value or values of (f1-f2)/2 ?

OpenStudy (anonymous):

335 and 351

OpenStudy (jamesj):

How many zeros does the function cos(2pi.t) have per second? 2.

OpenStudy (jamesj):

How many zeros does the function cos(4pi.t) have per second? 4.

OpenStudy (anonymous):

2

OpenStudy (jamesj):

Hence you need \[ \frac{f_1 - f_2}{2} = 4\pi \] Now solve. And notice that cos is an even function so the change in frequency can be plus or minus.

OpenStudy (anonymous):

so its 343-f2/2=4pi f2=343-8pi?

OpenStudy (anonymous):

but that answer isnt even one of the opstions :/

OpenStudy (jamesj):

And sorry: we need to normalize the original frequencies. If you have a sound wave with a frequency of f_1, that doesn't mean the function is cos(f1.t). It means the function is \[ \cos(2\pi f_1 t) \] So, we have \[ \frac{ 2\pi |f_2 - f_1|}{2} = 4\pi \]

OpenStudy (jamesj):

Therefore in fact, \[ f_2 = f_1 \pm 4 \]

OpenStudy (anonymous):

which yields 343-f2=4 so u get 347 and 339 correct?

OpenStudy (jamesj):

Yes, \[ f_2 = 343 \pm 4 = 347, 339 \]

OpenStudy (jamesj):

347 or 339

OpenStudy (anonymous):

ok, i will look over the rules but i get what you mean

OpenStudy (anonymous):

lol thx for taking so much time with me

OpenStudy (jamesj):

It's worth playing with Wolfram to see what two sound waves of close frequency look like, and to convince yourself that the closer the sounds around, the less beats you hear per second.

OpenStudy (jamesj):

If you ever play music or tune an instrument, this is very useful feature.

OpenStudy (anonymous):

interesting. i took a look at the link u sent, i get what you mean. fact they had such a small difference they had fewer beats/s